OpenStudy (anonymous):

Hello, I would really appreciate some help with the following question. Please show steps. Thankyou ∫ (-5z)/(z²+5) dz

OpenStudy (therealmeeeee):

ok Alright, first consider the following system from the coefficients: 8x + 5y + 3 = 0 5x + 2y + 4 = 0 Solving gives x = -14/9 and y = 17/9. Thus the substitution we have to make is u = x - (-14/9) = x + 14/9 ====> x = u - 14/9 v = y - 17/9 =============> y = v + 17/9 dy/dx = v' / u' = dv/du dv/du = [8(u - 14/9) + 5(v + 17/9) + 3] / [5(u - 14/9) + 2(v + 17/9) + 4] dv/du = [8u - 112/9 + 5v + 85/9 + 3] / [5u - 70/9 + 2v + 34/9 + 4] dv/du = [8u - 112/9 + 5v + 85/9 + 3] / [5u - 70/9 + 2v + 34/9 + 4] dv/du = [8u + 5v] / [5u + 2v] v' = [8u + 5v] / [5u + 2v] q = v/u ==> u = v/q v = qu v' = q + uq' = q + (v/q)q' q + (v/q)q' = [8(v/q) + 5v] / [5(v/q) + 2v] q + (v/q)q' = [5q + 8] / [2q + 5] q^2 + vq' = [(q)(5q + 8)] / [2q + 5] vq' = [(q)(5q + 8)] / [2q + 5] - q^2 vq' = (8q - 2q^3) / (2q + 5) v(dq/dv) = (8q - 2q^3) / (2q + 5) v dq = (8q - 2q^3) / (2q + 5) dv (2q + 5) / (8q - 2q^3) dq = (1/v) dv Finally we can integrate this. The left one is partial fractions all the way: ∫ (2q + 5) / (8q - 2q^3) dq = ∫ (1/v) dv (1/16) [-9 ln(2 - q) + 10 ln(q) - ln(q + 2)] = ln(v) + C (1/16) [-9 ln(2 - v/u) + 10 ln(v/u) - ln(v/u + 2)] = ln(v) + C v/u = (y - 17/9) / (x + 14/9) (1/16) [-9 ln((18 x - 9 y + 45)/(9 x + 14)) + 10 ln((y - 17/9) / (x + 14/9)) - ln((18 x + 9 y + 11)/(9 x + 14))] = ln(y - 17/9 ) + C

OpenStudy (anonymous):

I'm not getting it. Could you plz put our reply more in the form of my question?